OpenStudy (anonymous):

A player scored 94points in a single professional basketball game. He made a total of 60 baskets, consisting of free throws (worth two points) and foul shots (worth one point). Find the number of free throws and the number of foul shots that the player made

OpenStudy (anonymous):

\[x+y=60\]\[2x+y=94\] might be a good way to start

OpenStudy (anonymous):

F is free throws and f be the foul shots then total shots is total baskets = > F+f = 60 and total points is 2F+f = 94. => F = 34 and f = 26